# ZMP-HH-25/5

### v1

@@ 3.3 Uniqueness of the Emergent String
Before turning to the worldsheet theory on the lightest string, we address an important consistency check of the emergent string limit. Namely, while so far we have assumed that in the limit only a single string becomes tensionless at the leading rate, one can also engineer situations in which two (or even more) such strings become light equally fast. This is realised if instead of a growth sector as in (3.28), leading to only one tensionless string, there is least one more $k_{2}\in\mathcal{I}_{\rm II}$ such that in the asymptotic limit, $\bar{s}^{k_{1}}/\bar{s}^{k_{2}}=\text{const}$, whereas all other $\bar{s}^{k_{i}}$ still scale to infinity at a smaller rate. In particular, such behaviour arises at the intersection of two type II divisors where the singularity type does not enhance beyond type II. In this regime the tension of the string with charge $\mathbf{e}_{k_{2}}=(\delta^{ik_{2}})$ is asymptotically of the same order as the tension of the string with charge $\mathbf{e}_{k_{1}}$; furthermore both tensions scale as the square of the BPS particle tower mass identified in Section 3.1, and, importantly, this is the leading tower scale. [^24] If the singularity type enhances to type III or type IV, there arise additional BPS towers whose mass lies below the mass scale of the two strings associated with each of the type II divisors. In this case, the coexistence of two different such strings at the same energy is not problematic because none of these strings defines the new perturbative duality frame. One might hence conclude that there are two different emergent strings becoming weakly coupled at the same time. This would, however, be in conflict with the Emergent String Conjecture [8], according to which there can only be one emergent critical [^25] A way out would be if the two strings of the same asymptotic mass scale are not critical; however, in the Section 4 will be established for Tyurin degenerations irrespective of the details of the growth sector. string at a given point in moduli space such that this string defines the perturbative duality frame.
@@ 4 Type IIB/Heterotic Duality from the Worldsheet
From the perspective of the 10d Type IIB action, the spacetime can be viewed as a fibration of the Calabi–Yau threefold $V$ over the disk $\mathbf{D}$, which represents the two directions normal to the string in the four extended spacetime dimensions. Whereas the generic fibre $V_{z}$ of this fibration is smooth, the central fibre, $V_{0}$, realises the Tyurin degeneration described around (2.28) and depicted in Figure -957, i.e. $V_{0}=X_{1}\cup_{Z}X_{2}$. The condition for the resulting configuration to preserve four supercharges in 2d is then equivalent to the total space $\mathcal{V}$ of the fibration, as defined in (2.27), to be a non-compact Calabi–Yau fourfold.
@@ 4.1 The Worldsheet Spectrum
Additional bosonic zero modes arise from massless $p$-form fields of the 4d $\mathcal{N}=2$ supergravity that localise on the string at $z=0$. The components of such $p$-form fields transverse to the string locus give rise to scalar fields on the worldsheet. To find these, we must hence identify the $1$- and 2-form fields of the 4d $\mathcal{N}=2$ theory that localise to $\{z=0\}\subset\mathcal{V}$ [^28] Other $p$-form symmetries with $p>2$ would give rise to $(p-2)$-forms on the string worldsheet which are non-dynamical for $p>2$ in 2d.
@@
$B_{6}={\cal B}_{2}\wedge\tilde{\omega}_{\rm K3}\,,\qquad C_{6}=\tilde{\cal B}_{2}\wedge\tilde{\omega}_{\rm K3}\,,$ (4.12)
@@
$\mathbf{z}_{0}$ transverse coordinate $\in\mathbb{C}^{*}$ interacting $(2,2)$
@@
$\widetilde{\mathbf{b}}^{0}$ Reduction of $C_{6}$ over $H^{4}(Z)$ interacting $(1,1)$
@@
Total free $(20-\rho,3)$
interacting $(\rho+4,5)$
@@
To summarise, all zero modes on the string arising from massless higher-form fields of Type IIB supergravity are associated with harmonic forms on $Z$. In addition there are scalar modes arising from the deformation space of $Z$ inside $V_{0}$. If $V_{0}$ was smooth, there would be one complex deformation of the divisor $Z$ since $h^{2,0}(Z)=1$. However, $V_{0}$ is not smooth since it splits into two components $X_{1}$ and $X_{2}$ intersecting over $Z$. Geometrically, this degeneration corresponds to one of the directions normal to $Z$ shrinking to zero size. Hence, the deformation space of $Z$ has $\text{dim}_{\mathbb{R}}=1$, providing one real scalar on the string. [^32] In the examples of Tyurin degenerations that will be discussed in Sect. 5.3, one can geometrically understand the deformation space of $Z$ as follows. For the degeneration depicted in Fig. -955(b), it corresponds to the transverse direction of the cycles $\widehat{\gamma}_{1},\widehat{\gamma}_{2}$ inside the base space. In the degenration depicted in Fig. -954(b), it corresponds to the transverse direction of $\gamma$ inside the fibre. Together with the two modes transverse to the string in the extended 4d spacetime, in total we have eight right-moving scalars and 24 left-moving scalars. Since the worldsheet theory on the string preserves $\mathcal{N}=(0,4)$ supersymmetry, the right-moving scalars have fermionic partners to form complete multiplets of the superalgebra on the worldsheet. The left-and right-moving central charges of the 2d theory are hence
@@ 4.2 The Worldsheet Action
Kinetic term for $\mathbf{b}^{0}$ and $\widetilde{\mathbf{b}}^{0}$.
In addition to the zero modes arising from $C_{4}$, there are the modes along the string coming from $\mathcal{B}_{2}$ and ~$\tilde{\mathcal{B}}_{2}$, the 2-forms obtained by expanding $B_{6}$ and $C_{6}$ as described in equation (4.12). These 2-forms can be treated in analogy to the 2-forms $B^{i}_{2}$. Accordingly, the $z,\bar{z}$ components of $\mathcal{B}_{2}$ and ~$\tilde{\mathcal{B}}_{2}$ give real scalar modes along the string which we denote by $\mathbf{b}^{0}$ and $\widetilde{\mathbf{b}}^{0}$, respectively. From the kinetic term in 10d, we can define the function $\mathcal{S}$ in complete analogy to (4.15) and find
$\mathcal{S}(z-\mathbf{z}_{0},\bar{z}-\bar{\mathbf{z}_{0}},\mathbf{b}^{0},\widetilde{\mathbf{b}}^{0})=e^{2D(z,\bar{z})}\delta^{(2)}(z-\mathbf{z}_{0})\eta\int_{\mathbb{R}^{1,1}}\mathrm{d}^{2}\sigma\left[\partial_{\sigma}\mathbf{b}^{0}\partial^{\sigma}\mathbf{b}^{0}+\partial_{\sigma}\widetilde{\mathbf{b}}^{0}\partial^{\sigma}\widetilde{\mathbf{b}}^{0}\right]\,,$ (4.27)
@@
$S_{\rm WS,kin}(\mathbf{z}_{0},\bar{\mathbf{z}}_{0},\mathbf{b}^{0},\widetilde{\mathbf{b}}^{0})=\frac{1}{2}\int{\rm d}^{2}\sigma\,g_{00}(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})\left[\partial_{\sigma}\mathbf{b}^{0}\partial^{\sigma}\mathbf{b}^{0}+\partial_{\sigma}\widetilde{\mathbf{b}}^{0}\partial^{\sigma}\widetilde{\mathbf{b}}^{0}\right]\,,$ (4.29)
@@
such that the kinetic term for the fields $\mathbf{b}^{0}$ and $\widetilde{\mathbf{b}}^{0}$ depends on $\mathbf{z}_{0}$ and introduces interactions between these fields on the worldsheet.
@@
This leaves us with the kinetic term for the transverse modes to the string. On the one hand, these are the modes $(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})$ for which the kinetic term simply reads
$S_{\rm WS,kin}(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})=\int_{\mathbb{R}^{1,1}}\mathrm{d}^{2}\sigma\,e^{2D(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})}\,\partial_{\sigma}\mathbf{z}_{0}\partial^{\sigma}\bar{\mathbf{z}}_{0}\,.$ (4.31)
Thus, the kinetic term for the $\mathbf{z}_{0}$-modes is not a constant but, at the same time, does not depend on the other fields on the worldsheet. Finally, we have the single real field $\Phi$ describing the normal deformations of $Z$ in $V_{0}$. The modes of $\Phi$ completely decouple from the other modes on the string, since it is a modulus of the central fibre $V_{0}$ determining the intersection locus of the two (quasi-)Fano threefolds $X_{1}$ and $X_{2}$. Therefore, the kinetic term of $\Phi$ is simply
@@
i.e. $\Phi$ is a free field on the worldsheet.
@@ 4.3 Heterotic Dual
To analyse the target space of the string $\sigma$-model we start from the grouping of the string zero modes into free and interacting fields achieved in the previous section and summarised in Table 4.1. The number of free versus interacting fields depends on the embedding of the K3 surface $Z$ into the degenerating threefold $V_{0}$: There are $\text{dim}\,{\rm Gr}_{2}+1$ free fields $b^{a}$ and $\Phi$, where the scalars $b^{a}$ are associated with the 2-forms in $H^{2}(Z)$ that form the transcendental sub-lattice $\Lambda_{\rm trans}\subset\Gamma_{3,19}$ with signature $(2,b)$, see (2.40). All remaining modes in Table 4.1 are interacting. Of these, the interacting modes $b^{i}$ are associated with the polarization lattice $\Lambda_{\rm pol}$, the orthogonal complement of $\Lambda_{\rm trans}$ in $\Gamma_{3,19}$ induced by $H^{2}(X_{1})$ and $H^{2}(X_{2})$ on $H^{2}(Z)$, see (2.42), which has signature $(1,\rho)$, with $\rho=19-b$.
@@ 4.3.1 $20\geq\text{rk}(\Lambda_{\rm trans})\geq 4$ ($1\leq\rho\leq 17$)
We begin with the simplest case, corresponding to $1\leq\rho\leq 17$ (i.e. $2\leq b\leq 18$). In this situation $\Lambda_{\mathrm{trans}}$ contains two copies of the hyperbolic plane $U$, i.e. $U\oplus U\subseteq\Lambda_{\mathrm{trans}}$, such that the two right-moving scalars arising from $C_{4}$ are always paired with two left-moving scalars to give two real, free scalars on the worldsheet. These scalars are compact. Let us denote the two right(left)-moving scalars by $(b^{1}_{+(-)},b^{2}_{+(-)})$ and the real scalars by
@@
The real scalars $\mathbf{b}^{1,2}$ can be viewed as coordinates on the target space of the string $\sigma$-model. Furthermore, since they give rise to four ${\rm U}(1)$ gauge fields in the bulk, the corresponding target space factor must be a copy of $T^{2}$ with two of the Abelian gauge factors arising as KK-${\rm U}(1)$’s and the remaining two ${\rm U}(1)$’s coming from the winding of the string. More precisely, the winding and KK-${\rm U}(1)$’s of the two $S^{1}$ factors in $T^{2}$ can be identified with the fields $A^{a}_{\pm}$ as
@@
The remaining free scalars arising from $C_{4}$ are purely left-moving and realise a left-moving $U(1)^{17-\rho}$ algebra on the string worldsheet that gives rise to a $U(1)^{17-\rho}$ gauge group in the bulk. [^34] Subleading corrections to the leading order behaviour of the Hodge norm (4.25) would trigger (some of) the scalars $b^{a}$ to be interacting. On the heterotic side, this is associated with a perturbative breaking of the gauge group. On the other hand, the free real scalar $\Phi$ on the worldsheet does not lead to a gauge field in the bulk. The target space for $\Phi$ is thus non-compact $\mathbb{R}$.
@@
Similarly, the remaining left-moving scalar field $b^{i\neq 3}$ do not give rise to massless gauge fields in the bulk. If we turned off the interactions for these fields, these would give rise to ${\rm U}(1)$ gauge fields in the bulk. Turning on interactions for the worldsheet fields corresponds to Higgsing this Abelian gauge group in the bulk. The bulk hence contains a $U(1)^{\rho-1}$ gauge group that is completely Higgsed.
What remains to discuss are the fields $\mathbf{b}^{0}$ and $\widetilde{\mathbf{b}}^{0}$. Being compact real scalars, they can be viewed as coordinates on the target space of the $\sigma$-model and give rise to $S^{1}$’s that are non-trivially fibred over the space transverse to the string due to the non-trivial interaction of these fields with $\mathbf{z}_{0}$. Thus, the fields behave like the field $\mathbf{b}^{3}$ and together form a $T^{3}$. From the target space perspective, two of these three scalar fields form a $T^{2}$ that is non-trivially fibred over the $\mathbf{z}_{0}$-plane whereas the remaining coordinate can be viewed as the angular coordinate in the space transverse to the string. In the following, we view $(\mathbf{b}^{0},\mathbf{b}^{3})$ as the coordinates on the torus and $\widetilde{\mathbf{b}}^{0}$ as the angular coordinate in the extended space transverse to the string. Thus, the interaction between $\mathbf{z}_{0}$ and the fields $\mathbf{b}^{0}$ and $\mathbf{b}^{3}$ is manifest as a non-trivial fibration of the torus parametrised by $(\mathbf{b}^{0},\mathbf{b}^{3})$ over the $\mathbf{z}_{0}$-plane. On the other hand, the combination
$u\equiv\Phi\,e^{\widetilde{\mathbf{b}}^{0}}\,,$ (4.37)
of worldsheet fields parametrises the complex plane transverse to the string, i.e., $u\in\mathbb{C}$.
@@
$\mathcal{M}=T^{2}_{(\mathbf{b}^{1},\mathbf{b}^{2})}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{\widetilde{\mathbf{b}}^{0}}}\,,$ (4.38)
@@
$K3_{\rm het}=\left(T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}}\right);$ (4.39)
the gauge background breaks the original $G\times U(1)^{4}$ gauge group of the heterotic string to a rank $21-\rho$ group [^35] Recall that in this section we assume $\rho\geq 1$ such that the rank of the surviving gauge group is at most 20 which can hence be a subgroup of $E_{8}\times E_{8}\times U(1)^{4}$ (or $Spin(32)/\mathbb{Z}_{2}\times U(1)^{4}$). at a generic point on the Coulomb branch where the surviving gauge group is given by its Cartan subgroup. [^36] Since the $b^{i\neq 3}$ fields only interact with $\mathbf{z}_{0}$, the bundle breaking the heterotic gauge group has to be on the local K3.

### v2

@@ 3.3 Uniqueness of the Emergent String
Before turning to the worldsheet theory on the lightest string, we address an important consistency check of the emergent string limit. Namely, while so far we have assumed that in the limit only a single string becomes tensionless at the leading rate, one can also engineer situations in which two (or even more) such strings become light equally fast. This is realised if instead of a growth sector as in (3.28), leading to only one tensionless string, there is least one more $k_{2}\in\mathcal{I}_{\rm II}$ such that in the asymptotic limit, $\bar{s}^{k_{1}}/\bar{s}^{k_{2}}=\text{const}$, whereas all other $\bar{s}^{k_{i}}$ still scale to infinity at a smaller rate. In particular, such behaviour arises at the intersection of two type II divisors where the singularity type does not enhance beyond type II. In this regime the tension of the string with charge $\mathbf{e}_{k_{2}}=(\delta^{ik_{2}})$ is asymptotically of the same order as the tension of the string with charge $\mathbf{e}_{k_{1}}$; furthermore both tensions scale as the square of the BPS particle tower mass identified in Section 3.1, and, importantly, this is the leading tower scale. [^24] If the singularity type enhances to type III or type IV, there arise additional BPS towers whose mass lies below the mass scale of the two strings associated with each of the type II divisors. In this case, the coexistence of two different such strings at the same energy is not problematic because none of these strings defines the new perturbative duality frame. One might hence conclude that there are two different emergent strings becoming weakly coupled at the same time. This would, however, be in conflict with the Emergent String Conjecture [8], according to which there can only be one emergent critical [^25] A way out would be if the two strings of the same asymptotic mass scale are not critical; however, in Section 4 the criticality will be established for Tyurin degenerations irrespective of the details of the growth sector. string at a given point in moduli space such that this string defines the perturbative duality frame.
@@ 4 Type IIB/Heterotic Duality from the Worldsheet
From the perspective of the 10d Type IIB action, the spacetime can be viewed as a fibration of the Calabi–Yau threefold $V$ over the disk $\mathbf{D}$, which represents the two directions normal to the string in the four extended spacetime dimensions. Whereas the generic fibre $V_{z}$ of this fibration is smooth, the central fibre, $V_{0}$, realises the Tyurin degeneration described around (2.28) and depicted in Figure 3, i.e. $V_{0}=X_{1}\cup_{Z}X_{2}$. The condition for the resulting configuration to preserve four supercharges in 2d is then equivalent to the total space $\mathcal{V}$ of the fibration, as defined in (2.27), to be a non-compact Calabi–Yau fourfold.
@@ 4.1 The Worldsheet Spectrum
Additional bosonic zero modes arise from massless $p$-form fields of the 4d $\mathcal{N}=2$ supergravity that localise on the string at $z=0$. The components of such $p$-form fields transverse to the string locus give rise to scalar fields on the worldsheet. To find these, we must hence identify the $1$- and 2-form fields of the 4d $\mathcal{N}=2$ theory that localise to $\{z=0\}\subset\mathcal{V}$. [^28] Other $p$-form symmetries with $p>2$ would give rise to (p2)(p-2)-forms on the string worldsheet which are non-dynamical for $p>2$ in 2d.
@@
$B_{6}={\cal B}_{2}\wedge\tilde{\omega}_{\rm K3}\,,\qquad C_{6}={\cal C}_{2}\wedge\tilde{\omega}_{\rm K3}\,$, (4.12)
@@
$|\mathbf{z}_{0}|$ radial transverse coordinate $\in\mathbb{C}^{*}$ interacting $(1,1)$
$\arg(\mathbf{z}_{0})$ angular coordinate $\in\mathbb{C}^{*}$ free $(1,1)$
@@
$\mathbf{c}^{0}$ Reduction of $C_{6}$ over $H^{4}(Z)$ interacting $(1,1)$
@@
Total free $(21-\rho,4)$
interacting $(\rho+3,4)$
@@
To summarise, all zero modes on the string arising from massless higher-form fields of Type IIB supergravity are associated with harmonic forms on $Z$. In addition there are scalar modes arising from the deformation space of $Z$ inside $V_{0}$. If $V_{0}$ was smooth, there would be one complex deformation of the divisor $Z$ since $h^{2,0}(Z)=1$. However, $V_{0}$ is not smooth since it splits into two components $X_{1}$ and $X_{2}$ intersecting over $Z$. Geometrically, this degeneration corresponds to one of the directions normal to $Z$ shrinking to zero size. Hence, the deformation space of $Z$ has $\text{dim}_{\mathbb{R}}=1$, providing one real scalar on the string. [^32] In the examples of Tyurin degenerations that will be discussed in Sect. 5.3, one can geometrically understand the deformation space of $Z$ as follows. For the degeneration depicted in Fig. 5(b), it corresponds to the transverse direction of the cycles $\widehat{\gamma}_{1},\widehat{\gamma}_{2}$ inside the base space. In the degenration depicted in Fig. 6(b), it corresponds to the transverse direction of $\gamma$ inside the fibre. Together with the two modes transverse to the string in the extended 4d spacetime, in total we have eight right-moving scalars and 24 left-moving scalars. Since the worldsheet theory on the string preserves $\mathcal{N}=(0,4)$ supersymmetry, the right-moving scalars have fermionic partners to form complete multiplets of the superalgebra on the worldsheet. The left-and right-moving central charges of the 2d theory are hence
@@ 4.2 The Worldsheet Action
Kinetic term for $\mathbf{b}^{0}$ and $\mathbf{c}^{0}$.
In addition to the zero modes arising from $C_{4}$, there are the modes along the string coming from $\mathcal{B}_{2}$ and $\tilde{\mathcal{B}}_{2}$, the 2-forms obtained by expanding $B_{6}$ and $C_{6}$ as described in equation (4.12). These 2-forms can be treated in analogy to the 2-forms $B^{i}_{2}$. Accordingly, the $z,\bar{z}$ components of $\mathcal{B}_{2}$ and $\tilde{\mathcal{B}}_{2}$ give real scalar modes along the string which we denote by $\mathbf{b}^{0}$ and $\mathbf{c}^{0}$, respectively. From the kinetic term in 10d, we can define the function $\mathcal{S}$ in complete analogy to (4.15) and find
$\mathcal{S}(z-\mathbf{z}_{0},\bar{z}-\bar{\mathbf{z}}_{0},\mathbf{b}^{0},\mathbf{c}^{0})=e^{2D(z,\bar{z})}\delta^{(2)}(z-\mathbf{z}_{0})\eta\int_{\mathbb{R}^{1,1}}\mathrm{d}^{2}\sigma\left[\partial_{\sigma}\mathbf{b}^{0}\partial^{\sigma}\mathbf{b}^{0}+\partial_{\sigma}\mathbf{c}^{0}\partial^{\sigma}\mathbf{c}^{0}\right]\,$, (4.27)
@@
$S_{\rm WS,kin}(\mathbf{z}_{0},\bar{\mathbf{z}}_{0},\mathbf{b}^{0},\mathbf{c}^{0})=\frac{1}{2}\int{\rm d}^{2}\sigma\,g_{00}(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})\left[\partial_{\sigma}\mathbf{b}^{0}\partial^{\sigma}\mathbf{b}^{0}+\partial_{\sigma}\mathbf{c}^{0}\partial^{\sigma}\mathbf{c}^{0}\right]\,$, (4.29)
@@
such that the kinetic term for the fields $\mathbf{b}^{0}$ and $\mathbf{c}^{0}$ depends on $\mathbf{z}_{0}$ and introduces interactions between these fields on the worldsheet.
@@
This leaves us with the kinetic term for the transverse modes to the string. On the one hand, we have the single real field $\Phi$ describing the normal deformations of $Z$ in $V_{0}$. The modes of $\Phi$ completely decouple from the other modes on the string, since it is a modulus of the central fibre $V_{0}$ determining the intersection locus of the two (quasi-)Fano threefolds $X_{1}$ and $X_{2}$. Therefore, the kinetic term of $\Phi$ is simply
@@
i.e. $\Phi$ is a free field on the worldsheet. Finally, we have the modes corresponding to the complex field $(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})$. The worldsheet fields must arrange into full supermultiplets of the 2d $\mathcal{N}=(0,4)$ supersymmetry, and in particular the number of free right-moving scalars must come in multiples of four. [^34] By constrast, the purely right-moving supersymmetry algebra does not force the left-moving scalars within a hypermultiplet to be either all interacting or all free. This implies that the modes $(\mathbf{z}_{0},\bar{\mathbf{z}}_{0})$ must give a free and an interacting real scalar field. Since the kinetic terms for the interacting fields only depend on $|\mathbf{z}_{0}|$ we conclude that $|\mathbf{z}_{0}|$ is an interacting scalar whereas $\text{arg}(\mathbf{z}_{0})$ remains free, see (4.1) for a summary.
@@ 4.3 Heterotic Dual
To analyse the target space of the string $\sigma$-model we start from the grouping of the string zero modes into free and interacting fields achieved in the previous section and summarised in Table 4.1. The number of free versus interacting fields depends on the embedding of the K3 surface $Z$ into the degenerating threefold $V_{0}$: There are the free fields $b^{a}$, $\Phi$, and ${\rm arg}({\bf z}_{0})$, where the scalars $b^{a}$ are associated with the 2-forms in $H^{2}(Z)$ that form the transcendental sub-lattice $\Lambda_{\rm trans}\subset\Gamma_{3,19}$ with signature $(2,b)$, see (2.40). All remaining modes in Table 4.1 are interacting. Of these, the interacting modes $b^{i}$ are associated with the polarization lattice $\Lambda_{\rm pol}$, the orthogonal complement of $\Lambda_{\rm trans}$ in $\Gamma_{3,19}$ induced by $H^{2}(X_{1})$ and $H^{2}(X_{2})$ on $H^{2}(Z)$, see (2.42), which has signature $(1,\rho)$, with $\rho=19-b$.
@@ 4.3.1 $20\geq\text{rk}(\Lambda_{\rm trans})\geq 4$ ($1\leq\rho\leq 17$)
[…]
@@
The real scalars $\mathbf{b}^{1,2}$ can be viewed as coordinates on the target space of the string $\sigma$-model. Furthermore, since they give rise to four ${\rm U}(1)$ gauge fields in the bulk, the corresponding target space factor must be a copy of $T^{2}$ with two of the Abelian gauge factors arising as $K$-${\rm U}(1)$’s and the remaining two ${\rm U}(1)$’s coming from the winding of the string. More precisely, the winding and $K$-${\rm U}(1)$’s of the two $S^{1}$ factors in $T^{2}$ can be identified with the fields $A^{a}_{\pm}$ as
@@
The remaining free scalars arising from $C_{4}$ are purely left-moving and realise a left-moving $U(1)^{17-\rho}$ algebra on the string worldsheet that gives rise to a $U(1)^{17-\rho}$ gauge group in the bulk. [^35] Subleading corrections to the leading order behaviour of the Hodge norm (4.25) would trigger (some of) the scalars $b^{a}$ to be interacting. On the heterotic side, this is associated with a perturbative breaking of the gauge group. On the other hand, the free real scalars $\Phi$ and ${\rm arg}({\bf z}_{0})$ on the worldsheet do not lead to gauge fields in the bulk. The target space for $\Phi$ is non-compact $\mathbb{R}$, and for ${\rm arg}({\bf z}_{0})$ it is an angular coordinate that will be interpreted momentarily.
@@
Similarly, the remaining left-moving scalar fields $b^{i\neq 3}$ do not give rise to massless gauge fields in the bulk. If we turned off the interactions for these fields, these would give rise to ${\rm U}(1)$ gauge fields in the bulk. Turning on interactions for the worldsheet fields corresponds to Higgsing this Abelian gauge group in the bulk. The bulk hence contains a $U(1)^{\rho-1}$ gauge group that is completely Higgsed.
What remains to be discussed are the fields $\mathbf{b}^{0}$ and $\mathbf{c}^{0}$. Being compact real scalars, they can be viewed as coordinates on the target space of the $\sigma$-model and give rise to $S^{1}$’s that are non-trivially fibred over the space transverse to the string due to the non-trivial interaction of these fields with $\mathbf{z}_{0}$. Thus, the fields behave like the field $\mathbf{b}^{3}$ and together form a $T^{3}$. From the target space perspective, two of these three scalar fields form a $T^{2}$ that is non-trivially fibred over the $\mathbf{z}_{0}$-plane whereas the remaining coordinate can be viewed as the angular coordinate in the space transverse to the string. In the following, we view $(\mathbf{b}^{0},\mathbf{c}^{0})$ as the coordinates on the torus $T^{2}$. The interaction between $|\mathbf{z}_{0}|$ and the fields $\mathbf{b}^{0},\mathbf{c}^{0}$ and $\mathbf{b}^{3}$ is manifest as a non-trivial fibration of the torus parametrised by $(\mathbf{b}^{0},\mathbf{c}^{0})$ over the $|\mathbf{z}_{0}|e^{\mathbf{b}^{3}}$-plane. On the other hand, the combination
$u\equiv\Phi\,e^{\arg(\mathbf{z}_{0})}\,$, (4.36)
of worldsheet fields parametrises the complex plane transverse to the string, i.e., $u\in\mathbb{C}^{*}$.
@@
$\mathcal{M}=T^{2}_{(\mathbf{b}^{1},\mathbf{b}^{2})}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{c}^{0})}\to\mathbb{C}_{|\mathbf{z}_{0}|e^{{\mathbf{b}}^{3}}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{\arg(\mathbf{z}_{0})}}\,$, (4.37)
@@
$K3_{\rm het}=\left(T^{2}_{(\mathbf{b}^{0},\mathbf{c}^{0})}\to\mathbb{C}_{|\mathbf{z}_{0}|e^{{\mathbf{b}}^{3}}}\right)\,.$ (4.38)
The gauge background breaks the original $G\times U(1)^{4}$ gauge group of the heterotic string to a rank $21-\rho$ group [^36] Recall that in this section we assume $\rho\geq 1$ such that the rank of the surviving gauge group is at most 20 which can hence be a subgroup of $E_{8}\times E_{8}\times U(1)^{4}$ (or $Spin(32)/\mathbb{Z}_{2}\times U(1)^{4}$). at a generic point on the Coulomb branch where the surviving gauge group is given by its Cartan subgroup. [^37] Since the $b^{i\neq 3}$ fields only interact with $|\mathbf{z}_{0}|$, the bundle breaking the heterotic gauge group has to be on the local K3.